# Concatenated Steane code with single-flag syndrome checks

### v2

@@ III Noise models and fault-tolerant error correction decoders
A naive way to measure the eigenvalue of a stabilizer generator is to use a syndrome extraction circuit with a single, so-called bare ancilla, as shown in Fig. 2(a). However, in the circuit-level noise model, the number of faults that can be corrected by an error correction protocol with this kind of circuit might be less than the number of errors correctable by the code. This is because a single gate fault may cause a single-qubit error that can propagate throughout the protocol and become an error on the multiple data qubits. Here, we want to ensure that the EC protocol is fault tolerant; that is, whenever the total number of faults that occurred in the EC protocol is no more than the number of errors correctable by the underlying code, the EC protocol can correct the input errors as expected and does not cause the output errors which are not correctable by the next EC cycle  [^1] For the formal definition of FTEC, refer to the revised definition of FTEC in [32] which is extended from the definition in [21]; see also [2] for definition comparison.
@@ IV Finding the gate ordering for the concatenated Steane code
FIGURE 4: A CNOT ordering for the $[\![49,1,9]\!]$ code that ensures that the fault set $\mathcal{F}_{t}$ is distinguishable. Here, we only display the ordering of the weight-12 stabilizer generators.
@@
FIGURE 5: All possible $X$-type errors on the $[\![7,1,3]\!]$ Steane code with the syndrome $\vec{s}=(100)$ (where the order of the syndrome bits corresponds to red, green, and blue $Z$-type generators). Black and white vertices correspond to qubits with and without errors, respectively. When evaluating the weight parities on the same side of the triangle, errors with the same logical class always have the same weight parity.
@@
Here, we focus only on the first case where $E_{1}$ and $E_{2}$ are $Z$-type errors as the proof of the second case is similar. Because $E_{1}$ and $E_{2}$ are $Z$-type errors with the same syndrome, $E_{1}E_{2}$ is either a $Z$-type stabilizer or a logical $Z$ operator. Also, observe that $[L,E_{1}E_{2}]=LE_{1}E_{2}-E_{1}E_{2}L=\left((-1)^{b_{1}}-(-1)^{b_{2}}\right)E_{1}LE_{2}$ where $b_{x}=0$ if $E_{x}$ and $L$ commute and $b_{x}=1$ of they anticommute. Similarly, we have $\{L,E_{1}E_{2}\}=LE_{1}E_{2}+E_{1}E_{2}L=\left((-1)^{b_{1}}+(-1)^{b_{2}}\right)E_{1}LE_{2}$.
@@
FIGURE 6: A good CNOT ordering with effective distance 9 (a) and a bad CNOT ordering with effective distance 8 (b) for weight-12 generators of the $[\![49,1,9]\!]$ concatenated Steane code. The ordering starts with the red qubit, ends with the blue qubit, and the second qubit is the green qubit. The only difference between the two orderings is the swap of the red and the green qubits.
Consider possible $X$-type errors on the $[\![7,1,3]\!]$ Steane code with the same syndrome $\vec{s}=(100)$ depicted in Fig. 5 as an example (where the order of the syndrome bits corresponds to red, green, and blue $Z$-type generators). Here, we can see that two errors are logically equivalent if and only if their weight parities evaluated on the same side of the triangle are equal. This comes from the fact that the $Z$-type operator of weight 3 on each side of the triangle is a logical $Z$ operator of the $[\![7,1,3]\!]$ Steane code. That is, if we know exactly on which side of the triangle the weight parity of the error is being measured, we can find the logical class of the error accurately.
@@
FIGURE 7: (a) The circuit level performance of the $[\![49,1,9]\!]$ concatenated Steane code and the $[\![61,1,9]\!]$ 6.6.6 color code. The two codes are very close in performance, however, the pseudothreshold of the concatenated Steane code is roughly 14% higher than that of the 6.6.6 color code. The pseudothreshold is against the $p=2p/3$ line, denoted by the green dashed line on the graph. The dotted lines, which serve as guides to the eyes, have the same slope as $p^{5}$ and intersect at the pseudothresholds, showing that our protocol is distance-preserving. (b) Logical vs physical error rates under code capacity error model for the $[\![49,1,9]\!]$ concatenated Steane code and the $[\![61,1,9]\!]$ 6.6.6 color code. The pseudothreshold is against the $p=2p/3$ line, denoted by the green dashed line on the graph. The dotted lines have the same slope as $p^{5}$ and intersect at the pseudothresholds, showing that both error correcting codes have distance 9.
@@ V Numerical results
We generate the noisy circuit definitions in Python using Cirq [34], sample the circuits using Stim [35], and decode them with our C++ lookup table decoder [2]. Pseudothresholds are calculated based on the intersection with the $p_{L}=2p/3$ line of the linearly interpolated sample mean. For pseudothreshold errors, we calculate the intersections with the $p_{L}=2p/3$ line of the upper and lower bounds and use the one that is further from the pseudothreshold.
@@ VI Discussion
Note that the code distance is preserved in both codes and both noise models, so all fault combinations with up to 4 faults can be successfully corrected. Since the cases will 5 faults are more likely to happen than the cases with 6 or more faults, we conjecture that the logical error rate for each code and each model is mainly determined by the proportion of fault combinations with 5 faults that lead to decoding failure. The proportion varies with the noise model, so the ranking in the circuit-level noise model could be different from the ranking in the code capacity noise model. The reason that the $[\![61,1,9]\!]$ code performs worse in the circuit-level noise model could be because its protocol has more total locations on average compared to the $[\![49,1,9]\!]$ code in the same noise model, which leads to more possible decoding failure cases.
@@
It should be noted that the 4.8.8 color code of distance 9 [27] is also a $[\![49,1,9]\!]$ self-orthogonal CSS code, so it is natural to ask whether it performs better than the $[\![49,1,9]\!]$ concatenated Steane code in the circuit-level noise model. However, as shown in Appendix B, it is impossible to construct a flag FTEC protocol with only a single flag qubit per generator that preserves the code distance for the 4.8.8 color codes. Since more flags are required when decoding in the circuit-level noise model, the 4.8.8 color code seems worse than the concatenated Steane code in terms of qubit efficiency. We still do not know whether the 4.8.8 color code could perform better than the concatenated Steane code, but it seems unlikely since more flags would lead to more locations per round.
@@ VII Conclusion and outlook
To our knowledge, no self-orthogonal CSS code of distance 9 with less than 49 data qubits has been found yet. Thus, we believe that the level of error suppression that the $[\![49,1,9]\!]$ concatenated Steane code can achieve might be a promising target to demonstrate on an experimental platform where gate errors dominate over idling noise and at least 51 qubits with all-to-all qubit connectivity are available (assuming that fast measurement and reset on ancilla qubits are possible). Early fault tolerant demonstrations of Rydberg atom systems [17] and trapped ion systems [16, 36, 37, 38, 39], promise to have the scale as well as error characteristics to make them a promising target for our scheme. Our protocol would need to be significantly modified to work on platforms with high idling noise and constrained connectivity, for example, superconducting qubit architectures. We point out that the planar layout of the code allows for embedding it in a 2D architecture. However, for the weight-12 stabilizer generators, the single-flag scheme imposes degree 13 connectivity on the ancilla qubits, which is much higher than the currently available devices [14, 15]. It is possible to preserve planarity and low connectivity at the cost of an increased number of ancilla qubits. Dominant idling noise requires further optimizations to minimize the depth of our circuit and, hence, to reduce the number of locations where idling noise can occur. We leave this exploration to future work.
@@
VIII Acknowledgements
@@ Appendix A Explicit CNOT orderings
FIGURE 12: A weight-5 logical error on the 4.8.8 color code of distance 5 can be extended to a logical error of weight $5+2i$ on the code of distance $5+2i$ by adding $2i$ errors on the qubits that connect the old and the new red boundaries.
@@
In Step (2), we observe that any of the logical operators in the W2L or the W4L set can be extended to a logical operator of weight $5+2i$ on any 4.8.8 color code of distance $5+2i,i=1,2,3,...$. This is due to the fact that each logical operator has support on only one qubit on the red boundary (the base) of the code of distance 5. On the code of distance 7, an error of the same form can be connected to the red boundary by adding two data errors. We can repeat this process $i$ times for the code of distance $5+2i$. See Fig. 12 for an example.

### v3

@@ III Noise models and fault-tolerant error correction decoders
A naive way to measure the eigenvalue of a stabilizer generator is to use a syndrome extraction circuit with a single, so-called bare ancilla, as shown in Fig. 2(a). However, in the circuit-level noise model, the number of faults that can be corrected by an error correction protocol with this kind of circuit might be less than the number of errors correctable by the code. This is because a single gate fault may cause a single-qubit error that can propagate throughout the protocol and become an error on the multiple data qubits. Here we want to ensure that the EC protocol is fault tolerant according to the following definition from [31], which is extended from the definition of FTEC from [21] (see also [2] for comparison between definitions from [21] and [31]).
Definition 1.
Fault-tolerant error correction [31]
Let $t\leq\lfloor(d-1)/2\rfloor$ where $d$ is the distance of a stabilizer code. An error correction protocol is $t$-fault tolerant if the following two conditions are satisfied:
  1. 1.
    For any input codeword with an error that can arise from $r$ faults before the protocol and corresponds to the trivial cumulative flag vector, if $s$ faults occur during the protocol with $r+s\leq t$, ideally decoding the output state gives the same codeword as ideally decoding the input state.
  2. 2.
    If $s$ faults occur during the protocol with $s\leq t$, regardless of the number of faults that can cause the input error, the output state differs from any valid codeword by an error that can arise from $s$ faults and corresponds to the trivial cumulative flag vector.
In other words, whenever the total number of faults that occurred in the EC protocol is no more than the number of errors correctable by the underlying code, we want to ensure that the EC protocol can correct the input errors as expected and does not cause the output errors which are not correctable by the next EC cycle.
@@ IV Finding the gate ordering for the concatenated Steane code
FIGURE 4: A CNOT ordering for the $[\![49,1,9]\!]$ code that ensures that the fault set $\mathcal{F}_{t}$ is distinguishable. Here, we only display the ordering of the weight-12 stabilizer generators. The starting (1) and ending (12) qubits for each generator are highlighted with the same color.
@@
FIGURE 5: All possible $X$-type errors on the $[\![7,1,3]\!]$ Steane code with the syndrome $\vec{s}=(100)$ (where the order of the syndrome bits corresponds to red (top), green (bottom-right), and blue (bottom-left) $Z$-type generators). Black and white vertices correspond to qubits with and without errors, respectively. When evaluating the weight parities on the same side of the triangle, errors with the same logical class always have the same weight parity.
@@
FIGURE 6: A good CNOT ordering with effective distance 9 (a) and a bad CNOT ordering with effective distance 8 (b) for weight-12 generators of the $[\![49,1,9]\!]$ concatenated Steane code. The ordering starts with qubit 1 and 2, ends with qubit 12. The only difference between the two orderings is the swap of the first and second qubits.
@@
Here, we focus only on the first case where $E_{1}$ and $E_{2}$ are $Z$-type errors as the proof of the second case is similar. Because $E_{1}$ and $E_{2}$ are $Z$-type errors with the same syndrome, $E_{1}E_{2}$ is either a $Z$-type stabilizer or a logical $Z$ operator. Also, observe that $[L,E_{1}E_{2}]=LE_{1}E_{2}-E_{1}E_{2}L=\left((-1)^{b_{1}}-(-1)^{b_{2}}\right)E_{1}LE_{2}$ where $b_{x}=0$ if $E_{x}$ and $L$ commute and $b_{x}=1$ if they anticommute. Similarly, we have $\{L,E_{1}E_{2}\}=LE_{1}E_{2}+E_{1}E_{2}L=\left((-1)^{b_{1}}+(-1)^{b_{2}}\right)E_{1}LE_{2}$.
@@
Consider possible $X$-type errors on the $[\![7,1,3]\!]$ Steane code with the same syndrome $\vec{s}=(100)$ depicted in Fig. 5 as an example (where the order of the syndrome bits corresponds to red (top), green (bottom-right), and blue (bottom-left) $Z$-type generators). Here, we can see that two errors are logically equivalent if and only if their weight parities evaluated on the same side of the triangle are equal. This comes from the fact that the $Z$-type operator of weight 3 on each side of the triangle is a logical $Z$ operator of the $[\![7,1,3]\!]$ Steane code. That is, if we know exactly on which side of the triangle the weight parity of the error is being measured, we can find the logical class of the error accurately.
@@
FIGURE 7: (a) The circuit level performance of the $[\![49,1,9]\!]$ concatenated Steane code and the $[\![61,1,9]\!]$ 6.6.6 color code. The two codes are very close in performance, however, the pseudothreshold of the concatenated Steane code is roughly 14% higher than that of the 6.6.6 color code. The pseudothreshold is against the $p=2p/3$ line, denoted by the green dashed line on the graph. The dotted lines, which serve as guides to the eyes, have the same slope as $p^{5}$ and intersect at the pseudothresholds, showing that our protocol is distance-preserving. (b) Logical vs physical error rates under code capacity error model for the $[\![49,1,9]\!]$ concatenated Steane code and the $[\![61,1,9]\!]$ 6.6.6 color code. The pseudothreshold is against the $p=2p/3$ line, denoted by the black dash-dot line on the graph. The dotted lines have the same slope as $p^{5}$ and intersect at the pseudothresholds, showing that both error correcting codes have distance 9.
@@ V Numerical results
We generate the noisy circuit definitions in Python using Cirq [33], sample the circuits using Stim [34], and decode them with our C++ lookup table decoder [2]. Pseudothresholds are calculated based on the intersection with the $p_{L}=2p/3$ line of the linearly interpolated sample mean. For pseudothreshold errors, we calculate the intersections with the $p_{L}=2p/3$ line of the upper and lower bounds and use the one that is further from the pseudothreshold. We note that the $p_{L}=2p/3$ line represents the infidelity of any single-qubit pure state when it is sent through the depolarizing channel in which each single-qubit Pauli error occurs with probability $p/3$.
@@ VI Discussion
Note that the code distance is preserved in both codes and both noise models, so all fault combinations with up to 4 faults can be successfully corrected. Since the cases with 5 faults are more likely to happen than the cases with 6 or more faults, we conjecture that the logical error rate for each code and each model is mainly determined by the proportion of fault combinations with 5 faults that lead to decoding failure. The proportion varies with the noise model, so the ranking in the circuit-level noise model could be different from the ranking in the code capacity noise model. The reason that the $[\![61,1,9]\!]$ code performs worse in the circuit-level noise model could be because its protocol has more total locations on average compared to the $[\![49,1,9]\!]$ code in the same noise model, which leads to more possible decoding failure cases.
@@
It should be noted that the 4.8.8 color code of distance 9 [27] is also a $[\![49,1,9]\!]$ self-orthogonal CSS code, so it is natural to ask whether it performs better than the $[\![49,1,9]\!]$ concatenated Steane code in the circuit-level noise model. However, as shown in Appendix B, it is impossible to construct a flag FTEC protocol with only a single flag qubit per generator that preserves the code distance for the 4.8.8 color codes. Since more flags are required when decoding in the circuit-level noise model, the 4.8.8 color code seems worse than the concatenated Steane code in terms of qubit efficiency. We still do not know whether the 4.8.8 color code could perform better than the concatenated Steane code. It is possible that the 4.8.8 color code would perform worse since more flag qubits involved would lead to more locations per round. However, from our simulations it is clear that the situation is subtle and knowing the number of locations alone might not be enough to predict the performance.
@@ VII Conclusion and outlook
To our knowledge, no self-orthogonal CSS code of distance 9 with less than 49 data qubits has been found yet. Thus, we believe that the level of error suppression that the $[\![49,1,9]\!]$ concatenated Steane code can achieve might be a promising target to demonstrate on an experimental platform where gate errors dominate over idling noise and at least 51 qubits with all-to-all qubit connectivity are available (assuming that fast measurement and reset on ancilla qubits are possible). Early fault tolerant demonstrations of Rydberg atom systems [17] and trapped ion systems [16, 35, 36, 37, 38] promise to have the scale as well as error characteristics to make them a promising target for our scheme. We would like to emphasize that the performances of the concatenated Steane code and the 6.6.6 color code are very close under the considered noise model. This means that exposed to realistic noise, the ranking of the two codes might change. This issue might be especially pertinent on a 2D array of neutral atoms, where the concatenated Steane code might need more qubit movements than the 6.6.6 code. The more movements are expected due to the higher connectivity of the weight-12 stabilizer generators, and these moves might result in a significant increase in idling noise.
Our protocol would need to be significantly modified to work on platforms with high idling noise and constrained connectivity, for example, superconducting qubit architectures. We point out that the planar layout of the code allows for embedding it in a 2D architecture. However, for the weight-12 stabilizer generators, the single-flag scheme imposes degree 13 connectivity on the ancilla qubits, which is much higher than the currently available devices [14, 15]. It is possible to preserve planarity and low connectivity at the cost of an increased number of ancilla qubits. Dominant idling noise requires further optimizations to minimize the depth of our circuit and, hence, to reduce the number of locations where idling noise can occur. We leave this exploration to future work.
@@
VIII Data and code availability
All source code to reproduce the data and the actual data in this work are available to download [41].
IX Acknowledgements
@@ Appendix A Explicit CNOT orderings
FIGURE 12: A weight-5 logical error on the 4.8.8 color code of distance 5 can be extended to a logical error of weight $5+2i$ on the code of distance $5+2i$ by adding $2i$ errors on the qubits that connect the old and the new bottom boundaries (depicted by horizontal lines).
@@
In Step (2), we observe that any of the logical operators in the W2L or the W4L set can be extended to a logical operator of weight $5+2i$ on any 4.8.8 color code of distance $5+2i,i=1,2,3,...$. This is due to the fact that each logical operator has support on only one qubit on the bottom boundary (the base of the triangle) of the code of distance 5. On the code of distance 7, an error of the same form can be connected to the red boundary by adding two data errors. We can repeat this process $i$ times for the code of distance $5+2i$. See Fig. 12 for an example.